# Molecular De Novo Design through Transformer-based Reinforcement Learning

### 10 Jan 2024 v4

@@ Abstract
In this work, we introduce a method: REINVENT-Transformer to fine-tune a Transformer-based generative model for molecular de novo design. Leveraging the superior sequence learning capacity of Transformers over Recurrent Neural Networks (RNNs), our model can generate molecular structures with desired properties effectively. In contrast to the traditional RNN-based models, our proposed method exhibits superior performance in generating compounds predicted to be active against various biological targets, capturing long-term dependencies in the molecular structure sequence. The model’s efficacy is demonstrated across numerous tasks, including generating analogues to a query structure and producing compounds with particular attributes, outperforming the baseline RNN-based methods. Our approach can be used for scaffold hopping, library expansion starting from a single molecule, and generating compounds with high predicted activity against biological targets.
@@ Introduction
The vast expanse of chemical space, encompassing an order of magnitude from $10^{60}-10^{100}$ possible synthetically feasible molecules (Schneider & Fechner, 2005), presents formidable obstacles to drug discovery endeavors. In this colossal landscape, the task of pinpointing a molecule that simultaneously meets the prerequisites for bioactivity, drug metabolism and pharmacokinetic (DMPK) profile, and synthetic accessibility becomes an undertaking similar to the proverbial search for a needle in a haystack. Pioneering de novo design algorithms (Böhm, 1992; Gillet et al., 1994) have attempted to address this by employing virtual strategies to design and evaluate molecules, thereby condensing the vast chemical space into a more navigable realm for exploration.
Traditional de novo design models, based on Recurrent Neural Networks (RNNs), have proven effective in molecule generation tasks (Olivecrona et al., 2017). However, RNNs possess inherent architectural limitations, notably in their capability to capture long-term dependencies in sequential data, which can be particularly detrimental when modeling complex molecular structures. Recently, the Transformer architecture has emerged as a powerful alternative to RNNs in sequence modeling tasks across various domains. Some of the key advantages of Transformers over RNNs include:
@@
Early de novo design algorithms primarily focused on structure-based methods, where the aim was to develop ligands that precisely fit the binding pocket of a target, as highlighted in works by Böhm (Böhm, 1992) and Gillet et al. (Gillet et al., 1994). These methods, while effective in certain aspects, often resulted in molecules with suboptimal drug metabolism and pharmacokinetic (DMPK) properties, and posed challenges in synthetic tractability. In contrast, ligand-based approaches, which do not rely on the 3D structure of the target, were introduced to overcome some of these limitations. They involve creating a comprehensive virtual library of chemical structures, which are then evaluated using a scoring function (Ruddigkeit et al., 2013; Hartenfeller et al., 2012). However, as noted by Blundell et al. Blundell (1996), it is important to recognize that the effectiveness of ligand-based methods compared to structure-based ones is not definitive. Both approaches have their unique advantages and limitations, and the choice between them depends on the specific requirements and context of the drug design process.
Recently, generative models such as RNN-based methods have been used for de novo design of molecules (Segler et al., 2017; Gómez-Bombarelli et al., 2018; Yu et al., 2017). They have shown success in tasks like learning the underlying probability distribution over a large set of chemical structures, reducing the search over chemical space to only molecules seen as reasonable. Further fine-tuning of the models was done using reinforcement learning (RL) (Jaques et al., 2017), which showed considerable improvement over the initial model.
@@
Molecular assembly strategies, such as string-based approaches like SMILES and SELFIES (Weininger, 1988; Krenn et al., 2020), provide an efficient representation of molecules. Graph-based methods offer an intuitive two-dimensional representation of molecular structures, with nodes and edges representing atoms and bonds, respectively (Zhou et al., 2019; Jin et al., 2018). On the other hand, synthesis-based strategies aim to generate only synthesizable molecules, ensuring that the design aligns with real-world applications (Bradshaw et al., 2020; Gao et al., 2022a).
@@
Various optimization algorithms have been utilized for molecular design. Genetic Algorithms (GAs) mimic natural evolutionary processes and have been applied in molecule generation using both SMILES and SELFIES representations (Brown et al., 2019; Nigam et al., 2021). Bayesian optimization (BO) is another class of method that builds a surrogate for the objective function, with applications such as BOSS and ChemBO in the molecular domain (Moss et al., 2020; Korovina et al., 2020). Variational autoencoders (VAEs) offer a generative approach, mapping molecules to and from a latent space, with notable methods including SMILES-VAE and JT-VAE (Gómez-Bombarelli et al., 2018; Jin et al., 2018). Reinforcement Learning (RL) techniques, like REINVENT, have also been applied to tune models for molecule generation Olivecrona et al. (2017).
Furthermore, recent advancements in gradient ascent methods, such as Pasithea and Differentiable scaffolding tree (DST), have leveraged gradient-based optimization for molecular design (Shen et al., 2021; Fu et al., 2022).
@@
The evolution of molecular design methodologies has progressively addressed various challenges and limitations. The transition from RNN-based methods to more advanced generative models underscores a quest for improved handling of complex chemical structure representations and optimization. While RNNs brought significant progress, their inherent difficulty in capturing long-term dependencies in sequential data has been a notable shortcoming. This gap is precisely where the Transformer architecture, introduced by Vaswani et al. (Vaswani et al., 2023), brings its strengths to the fore. Its self-attention mechanism allows for a more nuanced and effective handling of sequence data, a critical aspect in molecular design where long-range interactions within molecules play a pivotal role.
@@
Parallelly, the field has seen the integration of reinforcement learning (RL) for fine-tuning generative models, as evidenced in the work by Jaques et al. (Jaques et al., 2017). RL’s ability to iteratively improve models based on a feedback loop aligns well with the demands of molecular design, where continuous refinement based on molecular properties is essential. The combination of RL with generative models has been shown to enhance the ability to navigate the vast chemical space more effectively, achieving better results in molecule generation.
@@
$m^{*}=\arg\max_{m\in\mathcal{M}}\mathcal{O}(m)$
Here, $m$ represents a molecular structure, while $\mathcal{M}$ is the expansive domain known as chemical space, encompassing all potential molecular candidates. The extent of $\mathcal{M}$ is overwhelmingly vast, for instance, around $10^{60}$ Bohacek et al. (1996). It’s presupposed that we have access to the actual value of a targeted property, symbolized by $\mathcal{O}(m):\mathcal{M}\rightarrow\mathcal{R}$, in which an oracle, $\mathcal{O}$, functions as an opaque mechanism. This oracle assesses specific chemical or biological attributes of a molecule $m$ and yields the real property $\mathcal{O}(m)$ as a scalar value. It’s important to note that the oracles do not provide an analytical form or derivatives of the properties. The most feasible oracles—either experimental procedures or high-fidelity simulations—often come with considerable expense. Therefore, an algorithm that can efficiently optimize the oracle within a feasible resource allocation is crucial. Such an algorithm would be a key component in the automation of molecular design, contributing significantly to advanced automated chemical design (ACD) Goldman et al. (2022) or function-driven autonomous synthesis Gao et al. (2022b).
@@
$\mu_{\mathbf{z}}=\frac{1}{k}\sum_{i=1}^{k}\mathbf{z}_{i},\quad\sigma_{\mathbf{z}}=\sqrt{\frac{1}{k}\sum_{i=1}^{k}\left(\mathbf{z}_{i}-\mu_{\mathbf{z}}\right)^{2}}$ (9)
@@
$J(\Theta)=-\sum_{t=1}^{T}\log P\left(x_{t}\mid x_{t-1},\ldots,x_{1}\right)$ (10)
@@
$NLL(S)=-\sum_{i=1}^{N}\ln P\left(X_{i}=T_{i}\mid X_{i-1}=T_{i-1}\ldots X_{1}=x_{1}\right)$ (11)
@@
$S(x)=\left[\prod_{i}p_{i}(x)^{w_{i}}\right]^{1/\sum_{i}w_{i}}$
$S(x)=\frac{\sum_{i}w_{i}*p_{i}(x)}{\sum_{i}w_{i}}$
@@
For any method that necessitates a database, we exclusively use the ZINC 250K dataset Irwin & Shoichet (2005). This dataset comprises approximately 250K molecules, selected from the ZINC database due to their pharmaceutical significance, manageable size, and widespread recognition. Both Screening Ahmed et al. (2018) and MolPAL Graff et al. (2021) conduct searches within this database. Additionally, generative models like VAEs Kingma & Welling (2014) and LSTMs Yu et al. (2019) are pretrained on it. Any fragments essential for JT-VAE (Jin et al., 2018), MIMOSA (Fu et al., 2021), and DST (Fu et al., 2022) are also derived from this very database.
@@
First, we compare two rule-based baselines, shown in the Table  1
@@
SMILES-LSTM HC
Iterative learning method using LSTM to understand molecular distribution in SMILES strings.
@@
Focuses on synthesizability of generated molecules.
@@
Our result is shown in Table. 2. From the table, we can observe that our method is better than the baseline method on multiple Oracles, which proves the effectiveness of the transformer in our problem.
@@
The evaluation results depict a thorough comparison between the REINVENT-Transformer (referred to as REINVENT-Trans) and other prominent models across multiple oracles. Randomly selected SMILES generated by different models can be seen in Table  4. And the corresponding chemical structures are shown in figure  2
@@
2. Graph-based Models: Both ‘Graph GA’ and ‘GP BO’ exhibit competitive performance in certain oracles like ‘Amlodipine_MPO’ and ‘Celecoxib_Rediscovery’ respectively. However, their performance isn’t consistently at the top across all oracles. This implies that while graph-based models can be effective in certain scenarios, they may not always generalize well across diverse tasks.
3. Genetic Algorithms: STONED (using SELFIES representation) achieves the highest score in the ‘Fexofenadine_MPO’ oracle. Genetic algorithms, despite their inherent stochasticity, have potential in some specific optimization tasks.
@@
We set a threshold=50 for the length of generated molecular string. If the generated string is longer than the threshold, it will be considered as "long", other it’s considered as "short" . From the Figure 4, we can see the our method REINVENT-Transformer has better average score when generating long sequences.
@@
We plotted the auc_topk curve and number of epoches is the x-axis. From the figure as follows, we can see that our method REINVENT-Transformer converges faster than REINVENT method.
From Fig. 5, the evolution of the average accuracy for the top 100 predictions is evident. Upon examination, across equivalent epochs, the mean accuracy of REINVENT-Transformer consistently surpasses that of REINVENT. This indicates a more expedient convergence rate for the REINVENT-Transformer compared to REINVENT.
Figure 5: Mean and Standard Deviation of avg_top100 over Epochs for REINVENT and REINVENT-Transformer on oracle Mestranol_Similarity
The avg_top100 curve initially displays a steep incline, eventually plateauing post approximately 6000 epochs. Notably, beginning from the 2500th epoch, the performance differential between REINVENT-Transformer and REINVENT significantly widens.
@@
Figure 6: Mean and Standard Deviation of auc_top100 over Epochs for REINVENT and REINVENT-Transformer on oracle Albuterol_Similarity
Furthermore, the AUC top100 curve for Albuterol Similarity is illustrated in Fig. 6. In this context, the differential in performance between REINVENT-Transformer and REINVENT is more nuanced. It isn’t until the 8000th epoch that a discernible gap emerges. Ultimately, the REINVENT-Transformer exhibits marginally superior performance relative to REINVENT in this scenario.
Figure 7: Mean and Standard Deviation of auc_top10 over Epochs for REINVENT and REINVENT-Transformer on oracle Mestranol_Similarity
In Fig. 7, the AUC top10 curve for Mestranol Similarity is presented. Contrasted with the average accuracy curve, this AUC curve demonstrates a milder inclination initially, followed by a pronounced rise. Specifically, for the REINVENT-Transformer, the mean AUC top10 consistently surpasses that of REINVENT. Although the disparity is subtle during the initial epochs, it becomes more pronounced post the 5000th epoch and remains so thereafter.

### 08 Mar 2024 v5

# REINVENT-Transformer: Molecular De Novo Design through Transformer-based Reinforcement Learning

@@ Abstract
In this work, we introduce a method: REINVENT-Transformer to fine-tune a Transformer-based generative model for molecular de novo design. Leveraging the superior sequence learning capacity of Transformers over Recurrent Neural Networks (RNNs), our model can generate molecular structures with desired properties effectively. In contrast to the traditional RNN-based models, our proposed method exhibits superior performance in generating compounds predicted to be active against various biological targets, capturing long-term dependencies in the molecular structure sequence. The model’s efficacy is demonstrated across numerous tasks, including generating analogues to a query structure and producing compounds with particular attributes, outperforming the baseline RNN-based methods. Our approach can be used for scaffold hopping, library expansion starting from a single molecule and generating compounds with high predicted activity against biological targets.
@@ Introduction
The vast expanse of chemical space, encompassing an order of magnitude from $10^{60}-10^{100}$ possible synthetically feasible molecules (Schneider & Fechner, 2005; Zhang et al., 2023), presents formidable obstacles to drug discovery endeavors. In this colossal landscape, the task of pinpointing a molecule that simultaneously meets the prerequisites for bioactivity, drug metabolism, and pharmacokinetic (DMPK) profile, and synthetic accessibility becomes an undertaking similar to the proverbial search for a needle in a haystack. Pioneering de novo design algorithms (Böhm, 1992; Gillet et al., 1994) have attempted to address this by employing virtual strategies to design and evaluate molecules, thereby condensing the vast chemical space into a more navigable realm for exploration.
Traditional de novo design models, based on Recurrent Neural Networks (RNNs), have proven effective in molecule generation tasks (Olivecrona et al., 2017; Gao et al., 2022a). However, RNNs possess inherent architectural limitations, notably in their capability to capture long-term dependencies in sequential data, which can be particularly detrimental when modeling complex molecular structures. Recently, the Transformer architecture has emerged as a powerful alternative to RNNs in sequence modeling tasks across various domains. Some of the key advantages of Transformers over RNNs include:
@@
Early de novo design algorithms primarily focused on structure-based methods, where the aim was to develop ligands that precisely fit the binding pocket of a target, as highlighted in works by Böhm (1992) and Gillet et al. (1994). These methods, while effective in certain aspects, often resulted in molecules with suboptimal drug metabolism and pharmacokinetic (DMPK) properties, and posed challenges in synthetic tractability. In contrast, ligand-based approaches, which do not rely on the 3D structure of the target, were introduced to overcome some of these limitations. They involve creating a comprehensive virtual library of chemical structures, which are then evaluated using a scoring function (Ruddigkeit et al., 2013; Hartenfeller et al., 2012). However, as noted by Blundell et al. Blundell (1996), it is important to recognize that the effectiveness of ligand-based methods compared to structure-based ones is not definitive. Both approaches have their unique advantages and limitations, and the choice between them depends on the specific requirements and context of the drug design process.
Recently, generative models such as RNN-based methods have been used for de novo design of molecules (Segler et al., 2017; Gómez-Bombarelli et al., 2018; Yu et al., 2017; Fu et al., 2022a). They have shown success in tasks like learning the underlying probability distribution over a large set of chemical structures, reducing the search over chemical space to only molecules seen as reasonable. Further fine-tuning of the models was done using reinforcement learning (RL) (Jaques et al., 2017), which showed considerable improvement over the initial model.
@@
Molecular assembly strategies, such as string-based approaches like SMILES and SELFIES (Weininger, 1988; Krenn et al., 2020), provide an efficient representation of molecules. Graph-based methods offer an intuitive two-dimensional representation of molecular structures, with nodes and edges representing atoms and bonds, respectively (Zhou et al., 2019; Jin et al., 2018). On the other hand, synthesis-based strategies aim to generate only synthesizable molecules, ensuring that the design aligns with real-world applications (Bradshaw et al., 2020; Gao et al., 2022b; Fu et al., 2022a; Chen et al., 2024).
@@
Various optimization algorithms have been utilized for molecular design. Genetic Algorithms (GAs) mimic natural evolutionary processes and have been applied in molecule generation using both SMILES and SELFIES representations (Brown et al., 2019; Nigam et al., 2021). Bayesian optimization (BO) is another class of method that builds a surrogate for the objective function, with applications such as BOSS and ChemBO in the molecular domain (Moss et al., 2020; Korovina et al., 2020). Variational autoencoders (VAEs) offer a generative approach, mapping molecules to and from a latent space, with notable methods including SMILES-VAE and JT-VAE (Gómez-Bombarelli et al., 2018; Jin et al., 2018). Reinforcement Learning (RL) techniques, like REINVENT, have also been applied to tune models for molecule generation Olivecrona et al. (2017). Furthermore, recent advancements in gradient ascent methods, such as Pasithea and Differentiable scaffolding tree (DST), have leveraged gradient-based optimization for molecular design (Shen et al., 2021; Fu et al., 2022b).
@@
The evolution of molecular design methodologies has progressively addressed various challenges and limitations. The transition from RNN-based methods to more advanced generative models underscores a quest for improved handling of complex chemical structure representations and optimization. While RNNs brought significant progress, their inherent difficulty in capturing long-term dependencies in sequential data has been a notable shortcoming. This gap is precisely where the Transformer architecture, introduced by Vaswani et al. (Vaswani et al., 2023), brings its strengths to the fore. Its self-attention mechanism allows for more nuanced and effective handling of sequence data, a critical aspect in molecular design where long-range interactions within molecules play a pivotal role.
@@
Parallelly, the field has seen the integration of reinforcement learning (RL) for fine-tuning generative models, as evidenced in the work by Jaques et al. (2017). RL’s ability to iteratively improve models based on a feedback loop aligns well with the demands of molecular design, where continuous refinement based on molecular properties is essential. The combination of RL with generative models has been shown to enhance the ability to navigate the vast chemical space more effectively, achieving better results in molecule generation (Olivecrona et al., 2017; Zhou et al., 2019; Fu et al., 2021a; Gao et al., 2022a; Wu et al., 2022).
@@
$m^{*}=\arg\max_{m\in\mathcal{M}}\mathcal{O}(m),$
Here, $m$ represents a molecular structure, while $\mathcal{M}$ is the expansive domain known as chemical space, encompassing all potential molecular candidates. The extent of $\mathcal{M}$ is overwhelmingly vast, for instance, around $10^{60}$ Bohacek et al. (1996). It’s presupposed that we have access to the actual value of a targeted property, symbolized by $\mathcal{O}(m):\mathcal{M}\rightarrow\mathcal{R}$, in which an oracle, $\mathcal{O}$, functions as an opaque mechanism. This oracle assesses specific chemical or biological attributes of a molecule $m$ and yields the real property $\mathcal{O}(m)$ as a scalar value. It is important to note that the oracles do not provide an analytical form or derivatives of the properties. The most feasible oracles—either experimental procedures or high-fidelity simulations—often come with considerable expense. Therefore, an algorithm that can efficiently optimize the oracle within a feasible resource allocation is crucial. Such an algorithm would be a key component in the automation of molecular design, contributing significantly to advanced automated chemical design (ACD) Goldman et al. (2022) or function-driven autonomous synthesis Gao et al. (2022c).
@@
$\mu_{\mathbf{z}}=\frac{1}{k}\sum_{i=1}^{k}\mathbf{z}_{i},\quad\sigma_{\mathbf{z}}=\sqrt{\frac{1}{k}\sum_{i=1}^{k}\left(\mathbf{z}_{i}-\mu_{\mathbf{z}}\right)^{2}}.$ (9)
@@
$J(\Theta)=-\sum_{t=1}^{T}\log P\left(x_{t}\mid x_{t-1},\ldots,x_{1}\right).$ (10)
@@
$NLL(S)=-\sum_{i=1}^{N}\ln P\left(X_{i}=T_{i}\mid X_{i-1}=T_{i-1}\ldots X_{1}=x_{1}\right).$ (11)
@@
$S(x)=\left[\prod_{i}p_{i}(x)^{w_{i}}\right]^{1/\sum_{i}w_{i}},$
$S(x)=\frac{\sum_{i}w_{i}*p_{i}(x)}{\sum_{i}w_{i}}.$
@@
For any method that necessitates a database, we exclusively use the ZINC 250K dataset Irwin & Shoichet (2005); Huang et al. (2021). This dataset comprises approximately 250K molecules, selected from the ZINC database due to their pharmaceutical significance, manageable size, and widespread recognition. Both Screening Ahmed et al. (2018) and MolPAL (Graff et al., 2021) conduct searches within this database. Additionally, generative models like VAEs (Kingma & Welling, 2014) and LSTMs (Yu et al., 2019) are pretrained on it. Any fragments essential for JT-VAE (Jin et al., 2018), MIMOSA (Fu et al., 2021b), and DST (Fu et al., 2022b) are also derived from this very database.
@@
First, we compare two rule-based baselines in Table 1.
@@
SMILES-LSTM HC Brown et al. (2019)
Iterative learning method using LSTM to understand the molecular distribution in SMILES strings.
@@
Focuses on the synthesizability of generated molecules.
@@
Our result is shown in Table. 2. From the table, we can observe that our method is better than the baseline method on multiple Oracles, which proves the effectiveness of the transformer in our problem. Our experiments mainly follow the benchmark paper Gao et al. (2022a).
@@
The evaluation results depict a thorough comparison between the REINVENT-Transformer (referred to as REINVENT-Trans) and other prominent models across multiple oracles. Randomly selected SMILES generated by different models can be seen in Table 4. And the corresponding chemical structures are shown in figure 2.
@@
2. Graph-based Models: Both ‘Graph GA’ and ‘GP BO’ exhibit competitive performance in certain oracles like ‘Amlodipine_MPO’ and ‘Celecoxib_Rediscovery’, respectively. However, their performance isn’t consistently at the top across all oracles. This implies that while graph-based models can be effective in certain scenarios, they may not always generalize well across diverse tasks.
3. Genetic Algorithms: STONED (using SELFIES representation) achieves the highest score in the ‘Fexofenadine_MPO’ oracle. Despite their inherent stochasticity, genetic algorithms have potential in specific optimization tasks.
@@
We set a threshold=50 for the length of the generated molecular string. If the generated string is longer than the threshold, it will be considered as "long", other it’s considered as "short". From Figure 4, we can see the our method REINVENT-Transformer has better average score when generating long sequences.
@@
We plotted the auc_topk curve and the number of oracle calls is the x-axis. From the figure as follows, we can see that our method REINVENT-Transformer converges faster than REINVENT method.
From Fig. 5, the evolution of the average accuracy for the top 100 predictions is evident. Upon examination, across equivalent number of oracle calls, the mean accuracy of REINVENT-Transformer consistently surpasses that of REINVENT. This indicates a more expedient convergence rate for the REINVENT-Transformer compared to REINVENT.
Figure 5: Mean and Standard Deviation of avg_top100 over oracle calls for REINVENT and REINVENT-Transformer on oracle Mestranol_Similarity
The avg_top100 curve initially displays a steep incline, eventually plateauing post approximately 6000 oracle calls. Notably, beginning from the 2500th oracle call, the performance differential between REINVENT-Transformer and REINVENT significantly widens.
@@
Figure 6: Mean and Standard Deviation of auc_top100 over oracle calls for REINVENT and REINVENT-Transformer on oracle Albuterol_Similarity
Furthermore, the AUC top100 curve for Albuterol Similarity is illustrated in Fig. 6. In this context, the differential in performance between REINVENT-Transformer and REINVENT is more nuanced. It isn’t until the 8000th oracle call that a discernible gap emerges. Ultimately, the REINVENT-Transformer exhibits marginally superior performance relative to REINVENT in this scenario.
Figure 7: Mean and Standard Deviation of auc_top10 over oracle calls for REINVENT and REINVENT-Transformer on oracle Mestranol_Similarity
In Fig. 7, the AUC top10 curve for Mestranol Similarity is presented. Contrasted with the average accuracy curve, this AUC curve demonstrates a milder inclination initially, followed by a pronounced rise. Specifically, for the REINVENT-Transformer, the mean AUC top10 consistently surpasses that of REINVENT. Although the disparity is subtle during the initial oracle calls, it becomes more pronounced post the 5000th oracle call and remains so thereafter.